A Complete Bibliography of Publications in IEEE Transactions on Privacy

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 1.00

Title word cross-reference

Cover [Ano24].

Editorial [Vai24].

Front [Ano24].

IEEE [Vai24]. Introduction [Vai24].

Privacy [Vai24].

Transactions [Vai24].

References

Anonymous:2024:FC

[Ano24] Anonymous. Front cover. IEEE Transactions on Privacy, 1(??):C1, 2024. CODEN ???? ISSN 2836-208X (electronic).

Vaidya:2024:EII

[Vai24] Jaideep Vaidya. Editorial introduction to the IEEE Transactions on Privacy. IEEE Transactions on Privacy, 1(??):1–2, 2024. CODEN ???? ISSN 2836-208X (electronic).